Hello! Could you please help me with the problem. It appears more complicated than I thought.
The trip from the house to the lake at 30mi/h takes 12 min longer then the return trip at 48mi/h. How far apart are the lake and the house.
Thank you

Answer by richwmiller(9135)

r*t=d
the distance is the same in both directions
let t equal the time from the lake
t+12/60=t+1/5=time for the trip to the lake
48t=30(t+1/5)
48t=30t+6
18t=6
t=1/3 hour
1/3*48=16 miles apart.